OpenStudy (anonymous):

find the following. assume that x can represent any real number. use absolute value when necessary.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{x^2-16x+64}\] \[\Large \sqrt{(x-8)^2}\] \[\Large |x-8|\] So \[\Large \sqrt{x^2-16x+64}= |x-8|\]